Colloquium

3:30 p.m. Tuesday, November 9, 2010 Boehm Hall 260

Grits and Group Theory at Your Local Cracker Barrel

Dr. Jennifer Schaefer

Assistant Professor of Mathematics Dickinson College

Abstract

In this talk, we will analyze a common game many of you have played at your local Cracker Barrel, namely, peg solitaire, using a little bit of basic group theory. Who would have known you could get a side of group theory with your grits?

3:00 p.m. refreshments served 3:30 p.m. talk begins

Deptartment of Mathematics Kutztown University of Pennsylvania Dr. Paul S. Ache III, Chair